OpenStudy (anonymous):

does the point (3,2) lie on the line 2x+5y=4? im guessing its yes(teacher told us) but im getting -2=0??

OpenStudy (amistre64):

when you input the values (x=3; y=2) into the equation; does it equal 4?

OpenStudy (amistre64):

2(3) +5(2) ?= 4 6 + 10 ?= 4 16 ?= 4 ...false

OpenStudy (anonymous):

well i get y=-2/5x-4/5 then i plug it in

OpenStudy (amistre64):

why go thru the hassle of rearranging an equation? the results are the same no matter how you do it; so far as you do it correctly that is :)

OpenStudy (amistre64):

2x+5y=4 y = (-2/5)x + 4/5

OpenStudy (anonymous):

hahah ok thank you!!! and so it does not lie on the line then?

OpenStudy (amistre64):

it does not lie on the line; unless you typoed it in the question

OpenStudy (anonymous):

ok ty!